OpenStudy (anonymous):

A kids’ splash pool and an adults’ pool are shaped like similar triangles. The dimensions of the kids’ pool are one-fifth of the dimensions of the adults’ pool. If the area of the adults’ pool is 225 square meters, what is the area of the kids’ pool? 112m^2 225 m^2 9 m^2 45 m^2 (I think its 9 m^2, but can you check it?)

OpenStudy (ganpat):

The dimensions of the kids’ pool are one-fifth of the dimensions of the adults’ So area for similar triangle should be of the ratio 1:5.. So A1/5 = 225/5 = 45... m2..

OpenStudy (anonymous):

For area, aren't you supposed to square the ratio?

OpenStudy (ganpat):

Area = 1/2 * base * ht... so say, b2 = b1/5 and h2 = h/5.. So Area 2 = 1/2 * b1/ 5 * b2/5 so A2 = A1/ 25 ... = 225/25 = 9 m2.. i guess, u r right... 9 shd be answer.. sorry, i was wrong..

OpenStudy (anonymous):

Thank you for reassuring my answer. (: